OpenStudy (anonymous):

If cosθ=−34 with π<θ<3π/2 , find the following: (a) cos (−θ ), (b) sec(−θ ), (c) sin(−θ ) . If cosθ=−34 with π<θ<3π/2 , find the following: (a) cos (−θ ), (b) sec(−θ ), (c) sin(−θ ) . @Mathematics

OpenStudy (mertsj):

No cosine can be smaller than -1 so this makes no sense.

OpenStudy (anonymous):

i meant cosθ=-3/4

OpenStudy (mertsj):

So this angle is in the third quadrant. If the cos is -3/4 then the side adjacent is 3 and the hypotenuse is 4 and the side opposite is sqrt7.So... sin = -sqrt7/4 tan = sqrt7/3 cot =3/sqrt7 (reciprocal of tan) sec = -4/3 (reciprocal of cos) csc = -4/sqrt7 (reciprocal of sin)

OpenStudy (mertsj):

But I doubt you related the problem correctly because they probably meant for it to be a 3-4-5 right triangle.

OpenStudy (anonymous):

this how the problem was represented by my professor

OpenStudy (anonymous):

then cos(-θ) BECOMES WHAT?